A total of Rs 12000 is invested in two funds paying 9% and 11% simple interest. If the interest for the first year is Rs 1180, How much of the Rs 12000, is invested at 9%?
Answer by mananth(16946)

Part I 9.00% per annum ------------- Amount invested =x
Part II 11.00% per annum ------------ Amount invested = y
12000
Interest----- 1180.00

Part I 9.00% per annum ---x
Part II 11.00% per annum ---y
Total investment
x + 1 y= 12000 -------------1
Interest on both investments
9.00% x + 11.00% y= 1180
Multiply by 100
9 x + 11 y= 118000.00 --------2
Multiply (1) by -9
we get
-9 x -9 y= -108000.00
Add this to (2)
0 x 2 y= 10000
divide by 2
y = 5000
Part I 9.00% Rs 7000
Part II 11.00% Rs. 5000